# On the Theoretical Limitations of Embedding-Based Retrieval

### v1

@@ 3.3 Consequences
Practically, this means:
  1. 1.
    For any fixed dimension $d$, there exists a binary relevance matrix which cannot be captured via $d$-dimensional embeddings (as there are matrices with arbitrarily high sign-rank). In other words, retrieval tasks whose qrel matrices have higher sign-rank are more difficult to capture exactly for embedding models, requiring higher embedding dimensions.
  2. 2.
    If we are able to embed a given matrix $A\in\{0,1\}^{m\times n}$ in a row-wise order-preserving manner in $d$ dimensions, this implies a bound on the sign rank of $2A-\mathbf{1}_{m\times n}$. In particular, this suggests a practical mechanism for determining an upper-bound on sign-rank for matrices via gradient descent optimization of free embedding representations.
4 Empirical Connection: Best Case Optimization
We have now established a theoretical limitation of embedding models based on the sign-rank of the qrel matrix and their embedding dimension $d$. Now we seek to show that this empirically as well.
To show the strongest optimization case possible, we design experiments where the vectors themselves are directly optimizable with gradient descent. [^5] This could also be viewed as an embedding model where each query/doc are a separate vector via a lookup table. We call this “free embedding” optimization, as the embeddings are free to be optimized and not constrained by natural language, which imposes constraints on any realistic embedding model. Thus, this shows whether it is feasible for any embedding model to solve this problem: if the free embedding optimization cannot solve the problem, real retrieval models will not be able to either. It is also worth noting that we do this by directly optimizing the embeddings over the target qrel matrix (test set). This will not generalize to a new dataset, but is done to show the highest performance that could possibly occur.
Experimental Settings
[…]
Results
Figure 2 shows that the curve fits a 3rd degree polynomial curve, with formula $y=-10.5322+4.0309d+0.0520d^{2}+0.0037d^{3}$ ($r^{2}$=0.999). Extrapolating this curve outward gives the critical-n values (for embedding size): 500k (512), 1.7m (768), 4m (1024), 107m (3072), 250m (4096). We note that this is the best case: a real embedding model cannot directly optimize the query and document vectors to match the test qrel matrix (and is constrained by factors such as “modeling natural language”). However, these numbers already show that for web-scale search, even the largest embedding dimensions with ideal test-set optimization are not enough to model all combinations.
5 Empirical Connection: Real-World Datasets
The free embedding experiments provide empirical evidence that our theoretical results hold true. However, they still are abstract - what does this mean for real embedding models? In this section we (1) draw connections from this theory to existing datasets and (2) create an trivially simple yet extremely difficult retrieval task for existing SOTA models.
5.1 Connection to Existing Datasets
Although it not possible to instantiate all combinations when using large-scale corpora, search evaluation datasets are a proxy for what any user would ask for and ideally would be designed to test many combinations, as users will do. In many cases, developers of new evaluations simply choose to use fewer queries due to cost or computational expense of evaluation. For example, QUEST’s query “Novels from 1849 or George Sand novels” combines two categories of novels with the “OR” operator – one could instantiate new queries to relate concepts through OR’ing other categories together. Similarly, with the rise of search agents, we see greater usage of hyper-specific queries: BrowseComp (Wei et al., 2025) has 5+ conditions per query, including range operators. With these tools, it is possible to sub-select any top-$k$ relevant set with the right operators if the documents are sufficiently expressive (i.e. non-trivial). Thus, that existing datasets choose to only instantiate some of these combinations is mainly for practical reasons and not because of a lack of existence.
In contrast to these previous works, we seek to build a dataset that evaluates all combinations of top-$k$ sets for a small number of documents. Rather than using difficult query operators like QUEST, BrowseComp, etc. (which are already difficult for reasons outside of the qrel matrix) we choose very simple query and documents to highlight the difficulty of representing all top-$k$ sets themselves.
5.2 The LIMIT Dataset
Dataset Construction
In order to have a natural language version of this dataset, we need some way to map combinations of documents into something that could be retrieved with a query. One simple way to do this is to create a synthetic version with latent variables for queries and documents and then instantiate it with natural language. For this mapping, we choose to use attributes that someone could like (i.e. Jon likes Hawaiian pizza, sports cars, etc. ) as they are plentiful and don’t present issues w.r.t. other items: one can like Hawaiian pizza but dislike pepperoni, all preferences are valid. We then enforce two constraints for realism: (1) users shouldn’t have too many attributes, thus keeping the documents short (less than 50 per user) and (2) each query should only ask for one item to keep the task simple (i.e. “who likes X”). We gather a list of attributes a person could like through prompting Gemini 2.5 Pro. We then clean it to a final 1850 items by iteratively asking it to remove duplicates/hypernyms, while also checking the top failures with BM25 to ensure no overlap.
Our last step is to choose a qrel matrix to instantiate these attributes. Although we could not prove the hardest qrel matrix definitively with theory (as the sign rank is notoriously hard to prove), we speculate based on intuition that our theoretical results imply that the more interconnected the qrel matrix is (e.g. dense with all combinations) the harder it would be for models to represent. [^8] See Appendix 10 for specific metrics that show the difference between LIMIT and other IR datasets. Following this, we use the qrel matrix with the highest number of documents for which all combinations would be just above 1000 queries for a top-$k$ of 2 (46 docs, since $\binom{46}{2}$ is 1035, the smallest above 1k).
Figure 4: Scores on the LIMIT small task (N=46) over embedding dimensions. Despite having just 46 documents, model struggle even with recall@10 and cannot solve the task even with recall@20.
Models
We evaluate the state-of-the-art embedding models including GritLM (Muennighoff et al., 2024), Qwen 3 Embeddings (Zhang et al., 2025), Promptriever (Weller et al., 2024b), Gemini Embeddings (Lee et al., 2025), Snowflake’s Arctic Embed Large v2.0 (Yu et al., 2024), and E5-Mistral Instruct (Wang et al., 2022, 2023). These models range in embedding dimension (1024 to 4096) as well as in training style (instruction-based, hard negative optimized, etc.). We also evaluate three non-single vector models to show the distinction: BM25 (Robertson et al., 1995; Lù, 2024), gte-ModernColBERT (Chaffin, 2025; Chaffin and Sourty, 2024), and a token-wise TF-IDF. [^9] This model turns each unique item into a token and then does TF-IDF. We build it to show that it gets 100% on all tasks (as it reverse engineers our dataset construction) and thus we do not include it in future charts.
We show results at the full embedding dimension and also with truncated embedding dimension (typically used with matryoshka learning, aka MRL (Kusupati et al., 2022)). For models not trained with MRL this will result in sub-par scores, thus, models trained with MRL are indicating with stars in the plots. However, as there are no LLMs with an embedding dimension smaller than 384, we include MRL for all models to small dimensions (32) to show the impact of embedding dimensionality.
Results
Figure 3 shows the results on the full LIMIT while Figure 4 shows the results on the small (46 document) version. The results are surprising - models severely struggle even though the task is trivially simple. For example, in the full setting models struggle to reach even 20% recall@100 and in the 46 document version models cannot solve the task even with recall@20.
We see that model performance depends crucially on the embedding dimensionality (better performance with bigger dimensions). Interestingly, models trained with more diverse instruction, such as Promptriever, perform better, perhaps because their training allows them to use more of their embedding dimensions (compared to models which are trained with MRL and on a smaller range of tasks that can perhaps be consolidated into a smaller embedding manifold).
For alternative architectures, GTE-ModernColBERT does significantly better than single-vector models (although still far from solving the task) while BM25 comes close to perfect scores. Both of these alterative architectures (sparse and multi-vector) offer various trade-offs, see section˜5.6 for analysis.
5.3 Is this Domain Shift?
Although our queries look similar to standard web search queries, we wondered whether there could be some domain shift causing the low performance. If so, we would expect that training on a training set of similar examples would significantly improve performance. On the other hand, if the task was intrinsically hard, training on the training set would provide little help whereas training on the test set would allow the model to overfit to those tokens (similar to the free parameterized experiments).
To test this we take an off the shelf embedding model and train it on either the training set (created synthetically using non-test set attributes) or the official test set of LIMIT. We use lightonai/modernbert-embed-large and fine-tune it on these splits, using the full dataset for in batch negatives (excluding positives) using SentenceTransformers (Reimers and Gurevych, 2019). We show a range of dimensions by projecting the hidden layer down to the specified size during training (rather than using MRL).
Results
Figure˜5 shows the model trained on the training set cannot solve the problem, although it does see very minor improvement from near zero recall@10 to up to 2.8 recall@10. The lack of performance gains when training in-domain indicate that poor performance is not due to domain shift. By training the model on the test set we see it can learn the task, overfitting on the tokens in the test queries. This aligns with our free embedding results, that it is possible to overfit to the $N=46$ version with only 12 dimensions. However, it is notable that the real embedding model with 64 dimensions still cannot completely solve the task, indicating that real world models are multiple times more limited than free-embeddings, exacerbating the limitations shown in Figure 2.
5.4 Effects of Qrel Patterns
As mentioned in previous sections, the crucial difference that makes LIMIT hard is that it tests models on more combinations of documents than typically used. Although this makes intuitive sense, here we ablate this decision and show that methods that do not test as many combinations (i.e. when the qrels are represented as a graph, have lower graph density) are easier empirically.
Experiment Setup
We instantiate LIMIT from four different qrel patterns: (1) random sampling from all combinations (2) a cycle-based setup where the next query is relevant to one document from the previous query and the following next document, (3) a disjoint pattern where each query is relevant to two new documents and (4) the pattern that maximizes the number of connections (n choose k) for the largest number of documents that fit in the query set (dense, our standard setup). For all configurations, we use the same setup as the main LIMIT (50k docs, 1k queries, kk=2, etc).
Results
We see in Figure 6 that all patterns except dense have relatively similar performance. However, moving to dense shows strikingly lower scores across the board for all models: GritLM drops 50 absolute recall@100, whereas E5-Mistral has an almost 10x reduction (40.4 vs 4.8 recall@100).
Figure 6: Model results from LIMIT datasets created with different qrel patterns. The dense qrel pattern that uses the maximum number of combinations is significantly harder than the other patterns. Note that the “dense” version is the main LIMIT shown in Figure 3.
5.5 Correlation with MTEB
@@
BEIR (used in MTEB v1) (Thakur et al., 2021; Muennighoff et al., 2022) has frequently been cited as something that embedding models have overfit to (Weller et al., 2025b; Thakur et al., 2025). We compare performance on LIMIT to BEIR in Figure 7. We see that performance is generally not correlated and that smaller models (like Arctic Embed) do worse on both, likely due to embedding dimension and pre-trained model knowledge.
5.6 Alternatives to Embedding Models
Our previous results show both theoretically and empirically that embedding models cannot represent all combinations of documents in their top-$k$ sets, making them unable to represent and solve some retrieval tasks. As current embedding models have grown larger (e.g. up to 4096), this has helped reduce negative effects for smaller dataset sizes. However, with enough combinations of top-$k$ sets the dimensionality would have to increase to an infeasible size for non-toy datasets.
Thus, our results show an interesting tradeoff: embeddings can represent a large amount of combinations but not all combinations. Although they are useful for first stage results to a degree, more expressive retriever architectures will be needed. We briefly discuss some of these below.
Cross-Encoders
Although not suitable for first stage retrieval at scale, they are already typically used to improve first stage results. However, is LIMIT challenging for rerankers also?
We evaluate a long context reranker, Gemini-2.5-Pro (Comanici et al., 2025) on the small setting as a comparison. We give Gemini all 46 documents and all 1000 queries at once, asking it to output the relevant documents for each query with one generation. We find that it can successfully solve (100%) all 1000 queries in one forward pass. This is in contrast to even the best embedding models with a recall@2 of less than 60% (Figure 4). Thus we can see that LIMIT is simple for state-of-the-art reranker models as they do not have the same limitations based on embedding dimension. However, they still have the limitation of being more computationally expensive than embedding models and thus cannot be used for first stage retrieval when there are large numbers of documents.
Multi-vector models
Multi-vector models are more expressive through the use of multiple vectors per sequence combined with the MaxSim operator (Khattab and Zaharia, 2020). These models show promise on the LIMIT dataset, with scores greatly above the single-vector models despite using a smaller backbone (ModernBERT, Warner et al. (2024)). However, these models are not generally used for instruction-following or reasoning-based tasks, leaving it an open question to how well multi-vector techniques will transfer to these more advanced tasks.
Sparse models
Sparse models (both lexical and neural versions) can be thought of as single vector models but with very high dimensionality. This dimensionality helps BM25 avoid the problems of the neural embedding models as seen in Figure 3. Since the $d$ of their vectors is high, they can scale to many more combinations than their dense vector counterparts. However, it is less clear how to apply sparse models to instruction-following and reasoning-based tasks where there is no lexical or even paraphrase-like overlap. We leave this direction to future work.
6 Conclusion
We introduce the LIMIT dataset, which highlights the fundamental limitations of embedding models. We provide a theoretical connection that shows that embedding models cannot represent all combinations of top-$k$ documents until they have a large enough embedding dimension $d$. We show these theoretical results hold empirically as well, through best case optimization of the vectors themselves. We then make a practical connection to existing state-of-the-art models by creating a simple natural language instantiation of the theory, called LIMIT, that these models cannot solve. Our results imply that the community should consider how instruction-based retrieval will impact retrievers, as there will be combinations of top-$k$ documents cannot represent.
Limitations
Although our experiments provide theoretical insight for the most common type of embedding model (single vector) they do not hold necessarily for other architectures, such as multi-vector models. Although we showed initial empirical results with non-single vector models, we leave it to future work to extend our theoretical connections to these settings.
We also did not show theoretical results for the setting where the user allows some mistakes, e.g. capturing only the majority of the combinations. We leave putting a bound on this scenario to future work and would invite the reader to examine works like Ben-David et al. (2002).
We have showed the theoretical connection that proves that some combinations cannot be represented by embedding models, however, we cannot prove apriori which types of combinations they will fail on. Thus, it is possible that there are some instruction-following or reasoning tasks they can solve perfectly, however, we do know that there exists some tasks that they will never be able to solve.
Acknowledgments
We thank Tanmaya Dabral, Zhongli Ding, Anthony Chen, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova for their helpful feedback.
[…]

### v2

@@ D.3 Consequences
The cyclotomic polynomial construction presented in Alon et al. (1985) implies that any qrel matrix has sign-rank at most $2k$, where $k$ is the largest number of documents for a particular query. The construction results in unnormalized vectors, however this can be easily adapted to normalized vectors by using one additional dimension. In agreement with Theorem˜1, this construction requires infinite precision in general, and is thus not feasible in practice.
D.4 Correlation with MTEB
@@
BEIR (used in MTEB v1) (Thakur et al., 2021; Muennighoff et al., 2022) has frequently been cited as something that embedding models have overfit to (Weller et al., 2025c; Thakur et al., 2025). We compare performance on LIMIT to BEIR in Figure 7. We see that performance is generally not correlated and that smaller models (like Arctic Embed) do worse on both, likely due to embedding dimension and pre-trained model knowledge.
Appendix E LLM Usage
LLMs were not used for any paper writing, only for coding help and title brainstorming.
Appendix F Metrics Measuring Qrel Graph Density
We show two metrics that treat the qrel matrix as a graph and show that LIMIT has unique properties compared to standard IR datasets (Table 2). We call these metrics Graph Density and Average Query Strength and describe them below.
Graph Density
We use the qrel matrix to construct the graph, where nodes are documents and an edge exists between two documents if they are both relevant to at least one common query.
For a given graph $G=(V,E)$ with $V$ being the set of nodes and $E$ being the set of edges, the graph density is defined as the ratio of the number of edges in the graph to the maximum possible number of edges. For an undirected graph, the maximum possible number of edges is $\frac{|V|(|V|-1)}{2}$. Thus, the density $\rho$ is calculated as:
$\rho=\frac{|E|}{\frac{|V|(|V|-1)}{2}}=\frac{2|E|}{|V|(|V|-1)}$
This metric indicates how connected the graph is; a density of 1 signifies a complete graph (all possible edges exist), while a density close to 0 indicates a sparse graph. For a qrel dataset, the
Average Query Strength
In a query-query graph where nodes are queries and edges represent similarity between queries (e.g., Jaccard similarity of their relevant documents), the strength of a query node $i$, denoted $s_{i}$, is defined as the sum of the weights of all edges incident to it. If $w_{ij}$ is the weight of the edge between query $i$ and query $j$, and $N(i)$ is the set of neighbors of query $i$, then the strength is:
$s_{i}=\sum_{j\in N(i)}w_{ij}$
The Average Query Strength $\bar{s}$ is the mean of these strengths across all query nodes in the graph:
$\bar{s}=\frac{1}{|V_{Q}|}\sum_{i\in V_{Q}}s_{i}$
where $V_{Q}$ is the set of all query nodes in the graph. This metric provides an overall measure of how strongly connected queries are to each other on average within the dataset, based on their shared relevant documents.
Comparisons to other datasets
We compare with standard IR Datasets such as NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and SciFact (Wadden et al., 2020). We also show an instruction-following dataset, FollowIR Core17 (Weller et al., 2024a). For all datasets, we use the test set only. The results in Table 2 show that LIMIT has significantly higher values for both of these metrics (i.e. 28 for query similarity compared to 0.6 or lower for the others).
Table 2: Metrics measuring the density of the qrel matrix. We see that LIMIT is significantly higher than other datasets, but that the closest are instruction-following datasets such as Core17 from FollowIR. Our empirical ablations suggest (although cannot definitively prove) that datasets with higher values here will be harder for retrieval models to represent.
Dataset Name Graph Density Average Query Strength
NQ 0 0
HotPotQA 0.000037 0.1104
SciFact 0.001449 0.4222
FollowIR Core17 0.025641 0.5912
LIMIT 0.085481 28.4653
Appendix G Table Forms of Figures
In this section we show the table form of various figures. For Figure 3 it is Table 5, Figure 4 in Table 4, Figure 2 in Table 6, and Figure 5 in Table 3.
Split Dim Recall@2 Recall@10 Recall@100
Test 32 85.5 98.4 100.0
Test 64 90.4 98.7 100.0
Test 128 93.1 99.5 99.9
Test 256 94.2 99.7 100.0
Test 384 95.6 99.6 100.0
Test 512 94.0 99.5 99.9
Test 768 96.1 99.8 100.0
Test 1024 96.5 99.8 100.0
Train 32 0.0 0.0 0.0
Train 64 0.1 0.3 2.2
Train 128 0.2 0.7 3.1
Train 256 0.0 0.0 0.4
Train 384 1.1 2.7 8.3
Train 512 0.7 2.3 9.8
Train 768 0.7 2.4 9.9
Train 1024 1.0 2.8 11.2
Table 3: Fine-tuning results in table form. See Figure 5 for the comparable plot.
Model Dim Recall@2 Recall@10 Recall@20
BM25 default 97.8 100.0 100.0
E5-Mistral 7B 32 7.9 32.6 56.2
E5-Mistral 7B 64 10.2 37.0 60.3
E5-Mistral 7B 128 14.5 41.9 65.9
E5-Mistral 7B 256 15.3 45.9 69.7
E5-Mistral 7B 512 22.2 54.7 74.8
E5-Mistral 7B 768 21.6 57.5 79.2
E5-Mistral 7B 1024 24.5 60.5 80.0
E5-Mistral 7B 2048 28.9 66.3 83.2
E5-Mistral 7B 3072 29.9 67.8 85.3
E5-Mistral 7B 4096 29.5 68.1 85.2
GTE-ModernColBERT default 83.5 97.6 99.1
GritLM 7B 32 7.8 33.5 56.3
GritLM 7B 64 9.4 35.9 59.6
GritLM 7B 128 14.2 42.7 64.9
GritLM 7B 256 17.3 46.2 68.3
GritLM 7B 512 21.8 55.6 76.7
GritLM 7B 768 23.8 58.1 80.1
GritLM 7B 1024 26.2 61.4 80.1
GritLM 7B 2048 33.0 69.1 86.2
GritLM 7B 3072 36.3 72.9 89.9
GritLM 7B 4096 38.4 75.4 90.5
Promptriever Llama3 8B 32 6.1 31.4 56.0
Promptriever Llama3 8B 64 8.9 35.8 62.3
Promptriever Llama3 8B 128 13.7 44.5 67.6
Promptriever Llama3 8B 256 18.5 52.1 74.1
Promptriever Llama3 8B 512 27.0 61.8 81.7
Promptriever Llama3 8B 768 35.5 69.0 84.7
Promptriever Llama3 8B 1024 38.0 73.5 89.1
Promptriever Llama3 8B 2048 46.2 83.6 94.2
Promptriever Llama3 8B 3072 49.2 87.3 96.6
Promptriever Llama3 8B 4096 54.3 90.0 97.7
Qwen3 Embed 32 8.3 30.6 53.9
Qwen3 Embed 64 9.4 35.5 57.6
Qwen3 Embed 128 11.6 38.3 60.8
Qwen3 Embed 256 14.3 41.6 63.8
Qwen3 Embed 512 16.1 43.7 66.0
Qwen3 Embed 768 17.2 45.3 69.3
Qwen3 Embed 1024 17.8 48.7 70.3
Qwen3 Embed 2048 19.5 51.5 72.4
Qwen3 Embed 3072 19.3 52.8 73.3
Qwen3 Embed 4096 19.0 52.3 73.8
Gemini Embed 2 4.2 23.0 45.5
Gemini Embed 4 4.2 21.9 46.0
Gemini Embed 8 4.9 23.2 47.0
Gemini Embed 16 5.2 24.7 47.5
Gemini Embed 32 6.3 25.2 50.6
Gemini Embed 64 6.9 30.6 55.0
Gemini Embed 128 7.7 37.0 62.9
Gemini Embed 256 14.6 46.9 69.7
Gemini Embed 512 23.3 58.4 77.9
Gemini Embed 768 28.8 67.5 84.5
Gemini Embed 1024 31.8 69.9 86.1
Gemini Embed 2048 31.9 70.3 87.1
Gemini Embed 3072 33.7 72.4 87.9
Snowflake Arctic L 32 8.3 30.3 53.8
Snowflake Arctic L 64 9.0 35.4 58.5
Snowflake Arctic L 128 12.7 41.3 65.1
Snowflake Arctic L 256 16.0 48.2 72.6
Snowflake Arctic L 512 16.7 51.3 74.1
Snowflake Arctic L 768 17.9 53.5 74.6
Snowflake Arctic L 1024 19.4 54.9 76.0
Snowflake Arctic L 2048 19.4 54.9 76.0
Snowflake Arctic L 3072 19.4 54.9 76.0
Snowflake Arctic L 4096 19.4 54.9 76.0
Table 4: Results for the LIMIT small version. See comparable Figure 4.
Model Dim Recall@2 Recall@10 Recall@100
E5-Mistral 7B 32 0.0 0.0 0.5
E5-Mistral 7B 64 0.0 0.1 0.4
E5-Mistral 7B 128 0.1 0.3 1.0
E5-Mistral 7B 256 0.4 0.9 1.9
E5-Mistral 7B 512 0.7 1.3 3.8
E5-Mistral 7B 768 0.9 1.7 4.3
E5-Mistral 7B 1024 0.9 1.8 5.9
E5-Mistral 7B 2048 1.0 1.9 6.8
E5-Mistral 7B 3072 1.3 2.0 7.7
E5-Mistral 7B 4096 1.3 2.2 8.3
Snowflake Arctic L 32 0.0 0.1 0.6
Snowflake Arctic L 64 0.2 0.4 1.7
Snowflake Arctic L 128 0.1 0.3 1.8
Snowflake Arctic L 256 0.2 0.8 2.5
Snowflake Arctic L 512 0.3 1.0 2.5
Snowflake Arctic L 768 0.4 1.1 3.1
Snowflake Arctic L 1024 0.4 0.8 3.3
Snowflake Arctic L 2048 0.4 0.8 3.3
Snowflake Arctic L 3072 0.4 0.8 3.3
Snowflake Arctic L 4096 0.4 0.8 3.3
GritLM 7B 32 0.0 0.0 0.8
GritLM 7B 64 0.0 0.1 0.3
GritLM 7B 128 0.1 0.3 1.3
GritLM 7B 256 0.1 0.4 2.8
GritLM 7B 512 0.6 1.8 6.5
GritLM 7B 768 1.5 3.1 8.7
GritLM 7B 1024 1.8 3.5 10.6
GritLM 7B 2048 2.3 4.3 11.8
GritLM 7B 3072 2.0 4.3 12.9
GritLM 7B 4096 2.4 4.1 12.9
Promptriever Llama3 8B 32 0.0 0.0 0.1
Promptriever Llama3 8B 64 0.0 0.0 0.3
Promptriever Llama3 8B 128 0.0 0.1 0.6
Promptriever Llama3 8B 256 0.2 0.4 1.8
Promptriever Llama3 8B 512 0.6 1.4 5.4
Promptriever Llama3 8B 768 1.3 3.1 8.7
Promptriever Llama3 8B 1024 2.1 4.4 12.8
Promptriever Llama3 8B 2048 3.2 6.5 18.1
Promptriever Llama3 8B 3072 2.9 6.3 17.8
Promptriever Llama3 8B 4096 3.0 6.8 18.9
Qwen3 Embed 32 0.0 0.1 1.1
Qwen3 Embed 64 0.0 0.2 1.0
Qwen3 Embed 128 0.3 0.4 1.8
Qwen3 Embed 256 0.4 0.8 3.2
Qwen3 Embed 512 0.6 1.3 3.3
Qwen3 Embed 768 0.7 1.5 3.8
Qwen3 Embed 1024 0.7 1.6 4.6
Qwen3 Embed 2048 0.9 1.7 4.7
Qwen3 Embed 3072 0.8 1.6 4.8
Qwen3 Embed 4096 0.8 1.8 4.8
Gemini Embed 2 0.0 0.0 0.1
Gemini Embed 4 0.0 0.0 0.0
Gemini Embed 8 0.0 0.0 0.0
Gemini Embed 16 0.0 0.0 0.0
Gemini Embed 32 0.0 0.0 0.0
Gemini Embed 64 0.0 0.0 0.3
Gemini Embed 128 0.0 0.1 0.3
Gemini Embed 256 0.0 0.1 1.2
Gemini Embed 512 0.2 1.1 3.6
Gemini Embed 768 0.9 2.5 7.6
Gemini Embed 1024 1.3 2.7 8.1
Gemini Embed 2048 1.5 3.1 8.5
Gemini Embed 3072 1.6 3.5 10.0
GTE-ModernColBERT default 23.1 34.6 54.8
BM25 default 85.7 90.4 93.6
Table 5: Results on LIMIT. See comparable Figure 3.
$d$ Critical-$n$
4 10
5 14
6 19
7 24
8 28
9 32
10 36
11 42
12 47
13 54
14 62
15 70
16 79
17 89
18 99
19 109
20 120
21 132
22 144
23 157
24 170
25 184
26 198
27 213
28 229
29 245
30 261
31 278
32 296
33 314
34 333
35 352
36 372
37 392
38 413
39 434
40 460
41 484
42 505
43 545
44 605
45 626
Table 6: Critical Values of n for different d values in the Free Embedding optimization experiments. See Figure 2 for the corresponding figure.
Model BEIR LIMIT R@100
Snowflake Arctic 55.22 3.3
Promptriever 56.40 18.9
E5-Mistral 57.07 8.3
GritLM 57.40 12.9
Gemini Embed 62.65 10.0
Qwen3 Embed 62.76 4.8
Table 7: BEIR vs LIMIT results. See Figure 7 for the comparable plot.